OpenStudy (ivyleaguebum):

Help :) cos(3x)=-1 Part 1) Solve this equation for 3x, finding all values on the interval [0,2π] that satisfy the equation. Part II) Use algebra to find all values of x between 0 and 2π that satisfy this equation.

OpenStudy (fibonaccichick666):

so how do you undo cosine?

OpenStudy (ivyleaguebum):

umm...would i have to put 3x by itself

OpenStudy (ivyleaguebum):

to solve for 3x

OpenStudy (fibonaccichick666):

yes, but first how do you undo cosine? just like how do you "undo" multiplication

OpenStudy (ivyleaguebum):

you undo multiplication with division, sooo cosine with secant?

OpenStudy (ivyleaguebum):

hellooo? I still kind of need help...anyone...^-^

OpenStudy (fibonaccichick666):

uhm not quite you use inverses

OpenStudy (ivyleaguebum):

ohh yeah, so cos^-1 or arccos

OpenStudy (fibonaccichick666):

yup to both sides

OpenStudy (fibonaccichick666):

for when you compute, so do the first step

OpenStudy (ivyleaguebum):

cos(3x)=-1 cos(3x)+1=0 is this correct

OpenStudy (fibonaccichick666):

I wouldn't start it that way because you cannot undo the cosine now

OpenStudy (fibonaccichick666):

your first step should be to get rid of the cosine

OpenStudy (ivyleaguebum):

okay...umm.. \[\cos(3x)/\cos^-1=-1/\cos^-1\]

OpenStudy (ivyleaguebum):

3x= -1/cos^-1

OpenStudy (ivyleaguebum):

That doesn't seem right...I dont know what im doing! aaaaaaaaa

OpenStudy (ivyleaguebum):

dont i have to set the equation to equal 0 or is that the case for this equation

OpenStudy (fibonaccichick666):

uhm, not quite, so cosine is a function like logarithms. ie lnx=1 e^lnx=e^1 x=e^1 or e^x=1 ln(e^x)=ln(1) x=0

OpenStudy (fibonaccichick666):

or what is f(f^-1(x))=?

OpenStudy (fibonaccichick666):

you can't divide by a function, you apply it's inverse to both sides to undo it

OpenStudy (ivyleaguebum):

im kind of going off of this equation 2cosx-sqrt2=0 at the end i got cosx= sqrt2/2 the found two solutions. i think thats what the question wants me to do

OpenStudy (fibonaccichick666):

huh? now i'm confused

OpenStudy (ivyleaguebum):

ok. 2cosx-sqrt 2=0 2cosx-sqrt2 + sqrt2= sqrt2 2cosx/2 = sqrt2/2 cosx=sqrt2/2, then i found pi/4 and 7pi/4 to be solutions

OpenStudy (ivyleaguebum):

but this equation is much more complicated in the sense that i cant just do those simple steps

OpenStudy (fibonaccichick666):

hmm

OpenStudy (fibonaccichick666):

on (0,pi)?

OpenStudy (fibonaccichick666):

(0,2pi)*

OpenStudy (ivyleaguebum):

yes.

OpenStudy (fibonaccichick666):

because I'm having a brain fart, do those equal 45 deg

OpenStudy (ivyleaguebum):

i plugged in this function in this website, but didnt too much help http://www.wolframalpha.com/input/?i=+cos%283x%29%3D-1+solve+

OpenStudy (fibonaccichick666):

I was helping you on that

OpenStudy (fibonaccichick666):

let's do a similar ex

OpenStudy (fibonaccichick666):

let's say we have sin(2u)=1

OpenStudy (fibonaccichick666):

my first step is to get rid of the sine

OpenStudy (ivyleaguebum):

okay, thats where im confused All the equations i have done with trig functions dont involve getting rid of cosine or sine and such

OpenStudy (ivyleaguebum):

so i dont know how to do that

OpenStudy (fibonaccichick666):

so i do this: \[sin^{-1}[sin(2u)]=sin^{-1}(1)\] \[\not{sin^{-1}[\not{sin}}(2u)]=sin^{-1}(1)\] \[(2u)=sin^{-1}(1)\]

OpenStudy (fibonaccichick666):

that's supposed to be cancelling the sin and sin^-1 since they equal 1

OpenStudy (ivyleaguebum):

isnt cos(3x)=-1 already simplfied for ex. If i would simplify 2cos(7x)+1=0 i would get cos(7x)=-1/2

OpenStudy (fibonaccichick666):

not yet

OpenStudy (fibonaccichick666):

if it's x= then it's simplified

OpenStudy (fibonaccichick666):

you want x by itself

OpenStudy (ivyleaguebum):

oh, i see how you did that

OpenStudy (ivyleaguebum):

i would get 3x= cos^-1(-1)

OpenStudy (fibonaccichick666):

so now, in mine we will solve for u

OpenStudy (fibonaccichick666):

but yes you will eventually get that. It's important to show your steps though

OpenStudy (ivyleaguebum):

i got pi/3 yes, i will

OpenStudy (ivyleaguebum):

since 3x = pi 3x/3=pi/3 x=pi/3

OpenStudy (ivyleaguebum):

cos^-1(-1)= pi thank for your help

OpenStudy (ivyleaguebum):

@_@

OpenStudy (fibonaccichick666):

i can buy that

OpenStudy (fibonaccichick666):

np

OpenStudy (fibonaccichick666):

medal?